# Time-VLM: Exploring Multimodal Vision-Language Models for Augmented Time Series Forecasting

### v1

@@ Abstract
Recent advancements in time series forecasting have explored augmenting models with text or vision modalities to improve accuracy. While text provides contextual understanding, it often lacks fine-grained temporal details. Conversely, vision captures intricate temporal patterns but lacks semantic context, limiting the complementary potential of these modalities. To address this, we propose Time-VLM, a novel multimodal framework that leverages pre-trained Vision-Language Models (VLMs) to bridge temporal, visual, and textual modalities for enhanced forecasting. Our framework comprises three key components: (1) a Retrieval-Augmented Learner, which extracts enriched temporal features through memory bank interactions; (2) a Vision-Augmented Learner, which encodes time series as informative images; and (3) a Text-Augmented Learner, which generates contextual textual descriptions. These components collaborate with frozen pre-trained VLMs to produce multimodal embeddings, which are then fused with temporal features for final prediction. Extensive experiments across diverse datasets demonstrate that Time-VLM achieves superior performance, particularly in few-shot and zero-shot scenarios, thereby establishing a new direction for multimodal time series forecasting.
@@ 1 Introduction
Time series forecasting plays a pivotal role across numerous domains, including finance (Idrees et al., 2019), climate (Karevan & Suykens, 2020), energy (Deb et al., 2017), and transportation (Zheng & Huang, 2020). Accurate predictions of future trends enable better decision-making, risk management, and resource allocation. Traditional statistical models, like ARIMA, have been the go-to methods for time series analysis, but they often struggle to capture complex, non-linear dependencies in large datasets. In contrast, deep learning approaches, e.g., RNNs (Medsker et al., 2001) and Transformer-based models (Li et al., 2019; Wu et al., 2021; Zhou et al., 2021; Liu et al., 2022a; Zhou et al., 2022; Nie et al., 2023), have revolutionized this field by learning intricate temporal patterns from raw data through techniques such as patch-based restructuring, auto-correlation mechanisms, and frequency decomposition strategies, delivering superior performance across various forecasting tasks. However, these models often struggle with domain generalization and scenarios with limited training data, particularly in few/zero-shot settings (Liang et al., 2024).
@@
  • Text-Augmented Models: Textual data, such as task-specific knowledge or domain events, provides valuable context for time series forecasting. For example, in financial forecasting, market descriptions or policy changes can enhance accuracy by adding insights beyond raw data. Approaches like Time-LLM (Jin et al., 2024) and UniTime (Liu et al., 2024b) map time series into textual representations, leveraging Large Language Models (LLMs) to capture contextual influences and infer predictions (Figure 1, right). However, these models face two challenges: (1) the modality gap between continuous time series and discrete text leads to information loss, and (2) pre-trained word embeddings are rarely optimized for time series forecasting, limiting fine-grained temporal pattern capture.
@@
  • Vision-Augmented Models: In contrast, transforming time series data into visual representations, such as through Gramian Angular Fields (GAF) or recurrence plots, enables models to identify and exploit underlying visual patterns, thereby facilitating the extraction of intricate temporal relationships using image-based feature learning techniques (e.g., CNNs) (Figure 1, left). Recent work (Wu et al., 2023b; Wang et al., 2024; Chen et al., 2024) demonstrates the natural alignment between time series and vision, as both are continuous and share structural similarities, allowing pretrained vision models to effectively extract hierarchical temporal features. However, these models often lack semantic context, limiting their ability to capture domain-specific knowledge or event-driven insights critical for real-world forecasting.
Despite advancements in text- and vision-based models, integrating both modalities with time series remains underexplored. Vision-Language Models (VLMs) (Radford et al., 2021) excel in aligning vision and text for tasks like multimodal reasoning. Building on this, we introduce Time-VLM, extending VLMs to include time series as a third modality. Each modality—text for semantic context, vision for spatial patterns, and time series for temporal dynamics—offers unique strengths. Current approaches often focus on single modalities, limiting their ability to harness the complementary benefits of all three. Our method leverages pretrained VLMs to effectively integrate temporal, visual, and textual data for augmented time series forecasting.
@@
By projecting time series into a unified vision-language semantic space, Time-VLM enables rich cross-modal interactions, combining the strengths of both modalities while mitigating their individual limitations. Specifically, Time-VLM introduces three key components: (1) a Retrieval-Augmented Learner that processes raw time series data through patch-based feature extraction and memory bank interactions to generate enriched temporal representations, capturing both local and global dependencies; (2) a Vision-Augmented Learner that adaptively transforms time series into images using multi-scale convolution, frequency encoding, and periodic encoding, preserving both fine-grained details and high-level structures; and (3) a Text-Augmented Learner that generates rich textual context (e.g., statistics and dataset descriptions) to complement the visual representations. These modules collaborate with VLMs to integrate temporal, visual, and textual modalities, producing accurate forecasts through a fine-tuned predictor.
Our key contributions are summarized as follows:
@@
  • We propose the first framework unifying temporal, visual, and textual modalities via pretrained VLMs, complementing their strengths for augmented time series forecasting.
@@
  • We introduce three components: a Retrieval-Augmented Learner for temporal feature enhancement via memory bank interactions, a Vision-Augmented Learner for adaptive time-series-to-image transformation, and a Text-Augmented Learner for contextual prompt generation.
  • Time-VLM achieves superior performance, particularly in few-shot and zero-shot scenarios, establishing a new direction for multimodal time series forecasting.
@@ 2 Related Work
The success of LLMs inspires their application to time series tasks. Methods like LLMTime (Gruver et al., 2023) and LLM4TS (Chang et al., 2023) tokenize time series data for autoregressive prediction but inherit LLMs’ limitations, such as poor arithmetic and recursive capabilities. Recent approaches, including GPT4TS (Zhou et al., 2023) and TimeLLM (Jin et al., 2024), project time series into textual representations to leverage LLMs’ reasoning abilities. However, they face challenges like the modality gap and lack of time series-optimized word embeddings, leading to potential information loss. UniTime (Liu et al., 2024b) and TimeFFM (Liu et al., 2024a) incorporate domain-specific instructions and federated learning, respectively, but remain constrained by their reliance on text alone.
@@
Vision emerges as a natural way to preserve temporal patterns. Early approaches use CNNs for matrix-formed time series (Li et al., 2020; Sood et al., 2021), while TimesNet (Wu et al., 2023b) introduces multi-periodic decomposition for unified 2D modeling. VisionTS (Chen et al., 2024) pioneers pre-trained visual encoders with grayscale time series images, and TimeMixer++ (Wang et al., 2024) advances the field with multi-scale frequency-based time-image transformations. Despite their effectiveness in temporal modeling, these methods often lack semantic context, hard to use high-level contextual information for prediction.
@@
VLMs like ViLT (Kim et al., 2021), CLIP (Radford et al., 2021), and ALIGN (Jia et al., 2021) transform multimodal understanding by aligning visual and textual representations. Recent advancements, like BLIP-2 (Li et al., 2023) and LLaVA (Liu et al., 2023), further enhance multimodal reasoning. However, VLMs remain underexplored for time series analysis. Our work bridges this gap by leveraging VLMs to integrate temporal, visual, and textual modalities, addressing the limitations of unimodal approaches.
@@ 3 Methodology
  • Retrieval-Augmented Learner (RAL): Extracts temporal features from raw time series patches, maintains a memory bank to interact with patch embeddings via multi-head self-attention and pooling mechanisms, preserving rich temporal feature representations and enhancing long-term dependency modeling for robust forecasting.
@@
  • Text-Augmented Learner (TAL): Generates contextual textual prompts for the input time series, including statistical features (e.g., mean, variance, trends), domain-specific context (e.g., electricity consumption patterns), and image descriptions. These prompts are encoded by a frozen VLM text encoder to produce textual embeddings.
@@ 3.1 Retrieval-Augmented Learner (RAL)
The RAL module extracts high-level temporal features via patch-based processing and memory-enhanced attention. It dynamically retrieves and integrates temporal patterns, adapting to complex time series structures to enhance forecasting performance. It operates in two key stages.
Patch Embedding: The input time series $x_{\text{enc}}\in\mathbb{R}^{B\times L\times D}$ (batch size $B$, sequence length $L$, variables $D$) is divided into overlapping patches of length patch_len with stride stride. Each patch is projected into a $d_{\text{model}}$-dimensional space via a linear layer, and positional embeddings are added to preserve temporal order. The resulting patch embeddings $E_{\text{patches}}\in\mathbb{R}^{B\times(D\cdot N_{\text{patches}})\times d_{\text{model}}}$ capture local temporal patterns, where $N_{\text{patches}}$ is the number of patches.
Memory-Enhanced Attention: A set of learnable memory queries $Q\in\mathbb{R}^{N_{\text{queries}}\times d_{\text{query}}}$ interact with patch embeddings through multi-head attention. Patch embeddings are projected into keys $K$ and values $V$, computing:
$\text{Memory}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{\text{query}}}}\right)V.$ (1)
To better capture complex temporal dynamics, we introduce a hierarchical memory structure with two levels:
  • Local Memory: Captures fine-grained patterns within individual patches for short-term dependencies.
  • Global Memory: Aggregates information across patches for long-range dependencies and high-level trends.
The two memories are fused via a gated mechanism:
$M_{\text{fused}}=\alpha\cdot M_{\text{local}}+(1-\alpha)\cdot M_{\text{global}},$ (2)
where $\alpha$ is a learnable gating parameter that adaptively balances the contribution of local and global memory.
The output $\mathbb{R}^{B\times N_{\text{queries}}\times d_{\text{query}}}$ stores high-level temporal patterns, enabling dynamic retrieval and fusion with other modalities. Each memory slot captures semantic and temporal relationships, allowing adaptive retrieval relevant patterns based on the input context. This hierarchical memory mechanism effectively addresses challenges in capturing long-range dependencies and complex temporal dynamics.
@@ 3.2 Vision-Augmented Learner (VAL)
The VAL module adaptively transforms time series $x_{\text{enc}}\in\mathbb{R}^{B\times L\times D}$ into images, preserving both fine-grained and high-level temporal patterns through following operations:
Frequency and Periodicity Encoding: To capture spectral and temporal dependencies, the VAL module applies two complementary encoding techniques to the input time series, explicitly add frequency and time-domain information.
@@
    Frequency Encoding: A Fast Fourier Transform (FFT) extracts frequency components as:
@@ 3.3 Text-Augmented Learner (TAL)
  • Range: minimum and maximum values,
  • Central tendency: median value,
  • Trend: upward/downward based on first-order differences,
  • Periodicity: dominant frequency components,
  • Task context: forecasting horizon and historical window,
  • Dataset characteristics: domain-specific descriptions.
These features are formatted into structured textual prompts. For example, the generated prompt might take the form:
“This is a time series transformed image highlighting its trends, periodic and change patterns. Dataset: {description}. Task: Forecast the next {pred_len} steps using the past {seq_len} steps. Input statistics: min value = {min_val}, max value = {max_val}, median value = {median_val}, the overall trend is {trend_direction}.”
When domain-specific knowledge is available (e.g., in medical diagnostics or financial analysis), TAL incorporates pre-defined textual descriptions. These are combined with dynamically generated prompts to enhance contextual understanding. The final textual representations are processed by the VLM text encoder, producing contextual embeddings that complement visual and temporal features.
By supporting both pre-defined and dynamically generated textual representations, TAL provides a flexible and adaptive mechanism for leveraging textual information in time series forecasting. This adaptability ensures that the model can effectively handle diverse scenarios, from general-purpose forecasting to domain-specific applications.
@@ 3.5 Optimization
The model is trained end-to-end using a mean squared error (MSE) loss to optimize forecasting performance. Given historical observations $\mathbf{X}\in\mathbb{R}^{N\times T}$ (with $N$ variables and $T$ time steps), the goal is to predict future values $\hat{\mathbf{Y}}\in\mathbb{R}^{N\times H}$ for $H$ time steps. The optimization objective is:
@@
where $\hat{\mathbf{Y}}_{h}$ and $\mathbf{Y}_{h}$ are predicted and ground truth values at time step $h$. The pre-trained VLM is frozen, and only lightweight components are optimized:
  • RAL: Patch embedding layer, memory queries, and attention mechanisms - capture temporal dependencies.
  • VAL: Multi-scale convolutional layers, frequency encoding, and periodicity encoding modules - transform time series into informative images.
  • Multimodal Fusion: Gating network, projection layers, and multi-head cross-attention - balance modality contributions (temporal, visual, and textual).
  • Predictor: fully connected layers - maps fused embeddings to forecasts.
@@ 4 Experiments
Datasets and Metrics. We evaluate Time-VLM on seven widely-used time series datasets spanning diverse domains, including energy consumption (ETTh1, ETTh2, ETTm1, ETTm2), weather forecasting, electricity load prediction (ECL, 321 variables), and traffic flow estimation (Traffic, 862 variables) (Zhou et al., 2021; Lai et al., 2018). These datasets, extensively adopted for benchmarking long-term forecasting models (Wu et al., 2023a), exhibit varying characteristics in sampling frequency, dimensionality, and temporal patterns. For short-term forecasting, we utilize the M4 benchmark (Makridakis et al., 2018), which encompasses marketing data at various sampling frequencies. Forecasting performance is evaluated using Mean Absolute Error (MAE) and Mean Squared Error (MSE), following standard practices in the field. Additional details on datasets and metrics are provided in Appendix A.1 and A.3.
@@
Implementation Details. We compare Time-VLM against state-of-the-art models using a unified evaluation pipeline, following the configurations in (Wu et al., 2023a) for fair comparison. ViLT (Kim et al., 2021) is the default backbone, with "vilt-b32-finetuned-coco". Other VLMs like CLIP and BLIP-2 are also supported. All models are trained with the Adam optimizer (learning rate $10^{-3}$, halved every epoch), a batch size of 32, and a maximum of 10 epochs with early stopping. Experiments are conducted on an Nvidia RTX A6000 GPU with 48GB memory. Additional optimization details are in Appendix A.2.

### v2

@@ Abstract
Recent advancements in time series forecasting have explored augmenting models with text or vision modalities to improve accuracy. While text provides contextual understanding, it often lacks fine-grained temporal details. Conversely, vision captures intricate temporal patterns but lacks semantic context, limiting the complementary potential of these modalities. To address this, we propose Time-VLM, a novel multimodal framework that leverages pre-trained Vision-Language Models (VLMs) to bridge temporal, visual, and textual modalities for enhanced forecasting. Our framework comprises three key components: (1) a Retrieval-Augmented Learner, which extracts enriched temporal features through memory bank interactions; (2) a Vision-Augmented Learner, which encodes time series as informative images; and (3) a Text-Augmented Learner, which generates contextual textual descriptions. These components collaborate with frozen pre-trained VLMs to produce multimodal embeddings, which are then fused with temporal features for final prediction. Extensive experiments demonstrate that Time-VLM achieves superior performance, particularly in few-shot and zero-shot scenarios, thereby establishing a new direction for multimodal time series forecasting. Code is available at https://github.com/CityMind-Lab/ICML25-TimeVLM.
@@ 1 Introduction
Time series data captures temporal signal evolution and underpins forecasting in diverse domains such as finance (Idrees et al., 2019), climate (Karevan & Suykens, 2020), energy (Deb et al., 2017), and transportation (Zheng & Huang, 2020). Accurate forecasting supports proactive risk mitigation, efficient resource allocation, and data-driven decision-making. Traditional models like ARIMA, while historically dominant, struggle to capture complex nonlinear patterns. In contrast, deep learning methods—from recurrent neural networks (RNNs) (Medsker et al., 2001) to Transformer-based architectures (Li et al., 2019; Wu et al., 2021; Zhou et al., 2021; Liu et al., 2022a; Zhou et al., 2022; Nie et al., 2023)—leverage innovations such as patch-based feature extraction, auto-correlation mechanisms, and frequency decomposition to model complex temporal dynamics. Despite their success, these models often fail to generalize across domains or adapt to data-limited scenarios, particularly few-shot and zero-shot settings (Liang et al., 2024).
@@
  • Text-Augmented Models: Textual data offers valuable semantics crucial for accurate forecasting. For instance, contextual descriptions or dataset statistics can significantly enrich the understanding of time series patterns (Figure 1, right). Methods like Time-LLM (Jin et al., 2024) and UniTime (Liu et al., 2024b) leverage the superior pre-trained inference capabilities of Large Language Models (LLMs) by mapping time series into textual representations. However, these approaches encounter two key challenges: (1) the modality gap between continuous time series and discrete text leads to information loss during representation alignment, and (2) pre-trained language knowledge is rare for capturing fine-grained temporal patterns, limiting their ability to learn nuanced dynamics.
@@
  • Vision-Augmented Models: Visual representations of time series—such as line graphs, recurrence plots, or grayscale images—enable models to exploit spatial patterns embedded within temporal data. By transforming sequences into images, methods such as CNNs, ViTs, and MAEs can extract hierarchical features that reveal latent temporal relationships (Figure 1, left). Recent studies (Wu et al., 2023b; Wang et al., 2024; Chen et al., 2024) have demonstrated the natural alignment between time series and vision, as both are continuous and share structural similarities, allowing pre-trained vision models to effectively encode temporal hierarchies. However, these methods lack semantic interpretability, restricting their capacity to incorporate domain-specific knowledge.
Despite advancements in text- and vision-augmented models, integrating both modalities with time series remains underexplored. Current approaches often focus on single modalities, failing to harness their combined strengths. To address this gap, we propose Time-VLM, a novel framework that leverages pre-trained VLMs to enhance time series forecasting by unifying temporal, visual, and textual information. VLMs provide a promising foundation, as they excel at aligning visual and textual modalities, making them well-suited for incorporating temporal information to unify the three modalities. By projecting time series into a unified vision-language semantic space, Time-VLM enables rich cross-modal interactions, combining the strengths of both modalities while mitigating their individual limitations. In this paradigm, each modality contributes uniquely: text provides semantic context, vision captures spatial-temporal patterns, and time series encodes sequential dynamics.
@@
Specifically, Time-VLM introduces three key components: (1) a Retrieval-Augmented Learner that processes raw time series data through patch-based feature extraction and memory bank interactions to generate enriched temporal representations, capturing both local and global dependencies; (2) a Vision-Augmented Learner that adaptively transforms time series into images using multi-scale convolution, frequency encoding, and periodic encoding, preserving both fine-grained details and high-level structures; and (3) a Text-Augmented Learner that generates rich textual context (e.g., statistics and dataset descriptions) to complement the visual representations. These modules collaborate with VLMs to integrate temporal, visual, and textual modalities, producing accurate forecasts through a fine-tuned predictor.
Our key contributions can be summarized as follows:
@@
  • We propose the first framework that unifies temporal, visual, and textual modalities by leveraging their complementary strengths for enhanced time series forecasting.
@@
  • We introduce a retrieval-augmented learner for hierarchical temporal feature enhancement, a vision-augmented learner for adaptive time-series-to-image transformation, and a text-augmented learner for contextual prompt generation, enabling seamless integration with VLMs.
  • Extensive evaluations show Time-VLM’s strong performance, especially under data-scarce conditions, offering a brand new paradigm for multimodal time series research.
@@ 2 Related Work
The success of LLMs has inspired their application to time series forecasting. Methods like LLMTime (Gruver et al., 2023) and LLM4TS (Chang et al., 2023) tokenize time series for autoregressive prediction but inherit limitations such as poor arithmetic and recursive reasoning. Approaches including GPT4TS (Zhou et al., 2023) and TimeLLM (Jin et al., 2024) project time series into textual representations to leverage LLMs’ reasoning capabilities, yet face challenges like the modality gap and limited temporal adaptability of word embeddings. UniTime (Liu et al., 2024b) and TimeFFM (Liu et al., 2024a) incorporate domain knowledge and federated learning, respectively, but remain constrained by their exclusive dependence on textual modeling.
@@
Vision provides a natural way to preserve temporal patterns in time series data. Early approaches apply CNNs to matrix-formed time series (Li et al., 2020; Sood et al., 2021), while TimesNet (Wu et al., 2023b) introduces multi-periodic decomposition for unified 2D modeling. VisionTS (Chen et al., 2024) pioneers the use of pre-trained visual encoders with grayscale time series images, and TimeMixer++ (Wang et al., 2024) advances the field through multi-scale frequency-based time-image transformations. Despite their effectiveness in temporal modeling, these methods often lack semantic context, limiting their ability to utilize high-level contextual information for prediction.
@@
VLMs like ViLT (Kim et al., 2021), CLIP (Radford et al., 2021), and ALIGN (Jia et al., 2021), have transformed multimodal understanding by aligning image and text representations. Recent progress, like BLIP-2 (Li et al., 2023) and LLaVA (Liu et al., 2023), further enhance multimodal reasoning. However, VLMs remain underexplored for time series analysis. Our work bridges this gap by leveraging VLMs to integrate temporal, visual, and textual modalities, addressing the limitations of unimodal approaches.
@@ 3 Methodology
  • Retrieval-Augmented Learner (RAL): Extracts temporal features from raw time series patches and maintains a memory bank to refine patch embeddings through multi-head self-attention and pooling mechanisms, thereby preserving rich temporal representations and enhancing long-term dependency modeling for robust forecasting.
@@
  • Text-Augmented Learner (TAL): Generates contextual textual prompts for input time series, including statistical features (e.g., mean, variance, trends), domain-specific context (e.g., electricity consumption patterns), and image descriptions. These prompts are encoded by a frozen VLM text encoder to produce textual embeddings.
@@ 3.1 Retrieval-Augmented Learner (RAL)
The RAL module extracts high-level temporal features via patch-based processing and retrieval-augmented memory mechanisms. It dynamically retrieves historical patterns and integrates them with current observations, adapting to complex time series structures. It operates in two key stages.
Patch Embedding: The input time series $x_{\text{enc}}\in\mathbb{R}^{B\times L\times D}$ is divided into overlapping patches of length $pl$ with stride $st$, where $B$, $L$, and $D$ denote the batch size, sequence length, and number of variables, respectively. Each patch is linearly projected into a $d_{\text{model}}$-dimensional latent space, and positional embeddings are added to preserve temporal order. This yields patch embeddings $E_{\text{p}}\in\mathbb{R}^{B\times N_{p}\times d_{\text{model}}}$, where $N_{p}=\frac{L-pl}{st}+1$ represents the total number of patches.
Retrieval-Augmented Memory: A memory bank with maximum capacity $M$ stores historical patch representations $\mathcal{M}\in\mathbb{R}^{M\times d_{\text{model}}}$. During each forward pass, the current patch embeddings are averaged across the temporal dimension and added to the memory bank using a circular buffer update strategy, ensuring that the most recent patterns are retained. To better capture temporal dynamics, we introduce a hierarchical memory structure:
  • Local Memory: Given current patch embeddings $P\in\mathbb{R}^{B\times N_{p}\times d_{\text{model}}}$, we retrieve top-$k$ similar patches from the memory bank $\mathcal{M}$ based on cosine similarity:
    $\text{sim}(P,\mathcal{M})=P\cdot\mathcal{M}^{\top},$ (1)
    where $\mathcal{M}\in\mathbb{R}^{M\times d_{\text{model}}}$ stores historical patch representations. The retrieved patches are processed through a two-layer MLP to extract local memory features:
    $M_{\text{local}}^{(i)}=\text{MLP}(\text{topk}(E_{p}^{(i)})),\quad i=1,\dots,B.$ (2)
    These features are averaged across patch dimension and combined with the original $P$ via a residual connection.
  • Global Memory: To capture long-range dependencies, we apply multi-head self-attention over the current patch embeddings $P$, yielding contextualized representations:
    $\text{Attn}(P)=\text{MultiHead}(Q,K,V),$ (3)
    where $Q,K,V$ are linear projections of $P$. The global memory is obtained by temporal averaging:
    $M_{\text{global}}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\text{Attn}(P)_{i}.$ (4)
The two memories are fused via element-wise addition:
$M_{\text{fused}}=M_{\text{local}}+M_{\text{global}},$ (5)
where $M_{\text{fused}}\in\mathbb{R}^{B\times N_{p}\times d_{\text{model}}}$ captures high-level temporal patterns. This fused representation supports dynamic retrieval and integration with other modalities (e.g., vision or text), enabling adaptive context-aware forecasting.
@@ 3.2 Vision-Augmented Learner (VAL)
The VAL module adaptively transforms the input time series $x_{\text{enc}}\in\mathbb{R}^{B\times L\times D}$ into image representations, enabling fine-grained and high-level temporal pattern extraction via the VLM vision encoder. The process is in three steps:
Frequency and Periodicity Encoding: To capture spectral and temporal dependencies, the VAL module applies two complementary encoding techniques to the input time series, explicitly adding frequency and time-domain information.
@@
    Frequency Encoding: A Fast Fourier Transform (FFT) extracts frequency components from raw input $x_{\text{enc}}$ as:
@@ 3.3 Text-Augmented Learner (TAL)
  • Statistical Properties: Value range (min/max), central tendency (median), and overall trend direction.
  • Contextual Information: Periodic description, task-specific parameters (input window length and forecasting horizon), and domain-specific dataset characteristics.
These features are formatted into structured textual prompts, such as the example shown in Figure 2(c). When domain-specific knowledge is available (e.g., in medical diagnostics or financial analysis), TAL incorporates pre-defined textual descriptions, which are combined with the dynamically generated prompts to enhance contextual understanding.
The final textual inputs are processed by the VLM text encoder, producing contextual embeddings that complement both visual and temporal features. This dual-path design supporting both static and dynamic textual ensures strong generalization across a wide range of applications, from generic forecasting tasks to specialized domain scenarios.
@@ 3.5 Optimization
The model is trained end-to-end using mean squared error (MSE). Given historical observations $\mathbf{X}\in\mathbb{R}^{N\times T}$ (with $N$ variables and $T$ time steps), the objective is to predict future values $\hat{\mathbf{Y}}\in\mathbb{R}^{N\times H}$ over $H$ steps:
@@
where $\hat{\mathbf{Y}}_{h}$ and $\mathbf{Y}_{h}$ denote predicted and ground-truth values at step $h$. The pre-trained VLM is kept frozen, and only lightweight components are optimized during fine-tuning:
  • RAL: Patch embedding, memory retrieval, and attention modules for temporal pattern learning;
  • VAL: Frequency/periodicity encoding and multi-scale CNNs for visual representation generation;
  • Prediction Head: A gate network and linear projection fuse multimodal features to generate final forecasts.
@@ 4 Experiments
Datasets and Metrics. We evaluate Time-VLM on seven widely used time series datasets across diverse domains: energy (ETTh1, ETTh2, ETTm1, ETTm2), weather, electricity (ECL), and traffic (Zhou et al., 2021; Lai et al., 2018). These datasets are commonly used for benchmarking long-term forecasting models (Wu et al., 2023a), and vary in frequency, dimensionality, and temporal characteristics. For short-term forecasting, we use the M4 benchmark (Makridakis et al., 2018), which includes marketing data across multiple frequencies. Performance is measured using Mean Absolute Error (MAE) and Mean Squared Error (MSE), following standard evaluation practices in this field. Additional details are provided in Appendices A.1 and A.3.
@@
Implementation Details. We compare Time-VLM against superior baselines using a unified evaluation pipeline under the same configurations as (Wu et al., 2023a) to ensure a fair comparison. ViLT (Kim et al., 2021) ("vilt-b32-finetuned-coco") serves as the default vision-language backbone; CLIP and BLIP-2 are also supported. All models are trained with Adam ($10^{-3}$ initial learning rate, halved per epoch), batch size 32, for up to 10 epochs with early stopping. Experiments run on Nvidia RTX A6000 GPU (48GB). More details are in Appendix A.2.